OpenStudy (anonymous):

Prove the relationship N1* N2 = HCF * LCM , for N1 = 125 & N2 =300

OpenStudy (welshfella):

find the prime factors 125 = 5*5*5 300 = 2*2*3*5*5 can you find the GCF and LCM from this info?

OpenStudy (anonymous):

ya

OpenStudy (welshfella):

ok the multiply them and see if it comes back to 300 *125 (= 37500)

OpenStudy (welshfella):

* then

OpenStudy (anonymous):

gcf = 25 lcm= 1500

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

1 medal ple

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

I hope I answered your quest :)